OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (-2, 5) and (6, -9). (4, -4) (2, -2) (1, -4)

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

thank you that helped

hartnn (hartnn):

welcome! if you need to verify your answer, just post what u got, i'll verify :)

OpenStudy (anonymous):

i got (2,-2)

hartnn (hartnn):

you are correct! good :)

OpenStudy (anonymous):

thank you im glad someone will help me

hartnn (hartnn):

you will find many users on OpenStudy who are dying to help you!! ;)

OpenStudy (anonymous):

ok thanks i just masde an account like a few days ago didnt know how good this sight was but i know now that its reliable